# Portable, Massively Parallel Implementation of a Material Point Method for Compressible Flows

### v1

@@ 3 Portable MPM Implementation
    Data generation: the creation of simulations input is performed separately from code execution. For simple geometries it is handled via ad-hoc C++ scripts. For complex geometries requiring a careful computation of the signed distance function, an Octave script is used instead, to take advantage of Octave built-in functions. In both cases, since objects are considered at rest, normal unit vectors are computed only once, and boundary cells are marked with a flag corresponding to the boundary condition. Finally, the initial data file is saved in JSON format.
@@
      Equations (6) and (8) are applied to map particles mass, momentum and force to the grid.
@@
      Velocity and acceleration G2P (7) is applied, and next, velocity is updated. Acceleration itself is not stored, but computed as the result coming from a local G2P of the relevant nodes ratio between force and mass. Following standard MPM [12], we store the G2P velocity in a temporary velocity vector $\mathbf{\overline{V}}$ and don’t use it to update particles velocity; however, we postpone its usage.
@@
      Particles strain rate tensor (velocity gradient) is computed as the G2PD (9) of the ratio between grid nodes momentum and mass, and then particles properties are updated.
@@
      Optionally, a re-ordering algorithm is applied to keep data locality, i.e. ti keep contiguous particles in the physical simulation also close in memory. The algorithm is based on an indexing vector built iterating over the structure that stores information about which particles are contained in every grid cell, such that the n-th element of the ordering vector indicates the current sorting particle index which has to be at n-th position after re-ordering.
@@
Algorithm 1 summarise the for loop procedure.
@@
$\qquad m=\mathrm{P2G}(M),\ m\mathbf{v}=\mathrm{P2G}(M\mathbf{V}),\ \mathbf{f}=\mathrm{P2GD}(P\tfrac{M}{\rho})$
@@
$\qquad\mathbf{V}^{\prime}=\mathbf{V}+\Delta t*\mathbf{A}$
@@
$\qquad E^{\prime}=E+\Delta t\dfrac{-P}{\rho}\nabla\cdot\mathbf{V}$
@@
This algorithm is of special interest for acceleration on GPUs since it can be written in terms of the four kernels (6), (7), (8), (9), which presents interesting parallelisation potentialities. Grid-to-Particles operations are intrinsically Single Instruction Multiple Data (SIMD) parallelizable, since data is projected from grid nodes to particles, and thus particles are independent from one another during writing stage. Particles-to-Grid kernel presents the same potentialities, however, if they are implemented strictly following their analytical definition, they present a data race: different particles need access in both read and write mode to the same grid nodes.
In this work, we chose to patch the P2G and G2P kernels written in the external quadgrid library to adapt them to GPU execution via Thrust rather than writing ad hoc versions in the code, at the expense of some further optimisation that could have been added by exploiting the peculiarities of our test cases, to arrive at more general and portable conclusions. In any case, preliminary tests have shown that applying a kernel-fusion approach [38] where possible would not have affected the execution time order of magnitude.
A P2G implementation is then based on the following pillars:
  1. 1.
    The definition of a functor for each kernel. For example, the initial P2G on mass could be implemented as follows: [^4] To compile CUDA or Thrust device code decorators are necessary; for example, the functor could be modified as: __device__ void operator () (idx_t k), and the __device__ macro could be defined as empty during compilation for portability, in case a non-CUDA aware compiler is used for different architectures.
    1struct functor_t
    2{
    3 real_t hx, hy;
    4 real_iter_t x, y, M, m;
    5 idx_iter_t ptcl_to_grid;
    6 functor_t (){};
    7 void operator () (idx_t k)
@@ A P2G implementation is then based on the following pillars:
    9 real_t N = 0.0, xx = x[k], yy = y[k];
    10 idx_t row = grid_to_row (ptcl_to_grid[k]);
    11 idx_t col = grid_to_col (ptcl_to_grid[k]);
@@
    14 N = shape_function (xx, yy, inode, col, row, hx, hy);
    15 idx_t i = global_node(inode, col, row);
    16 m[i] += N * M[k];
    17 }
    18 }
    19};
  2. 2.
    The definition of the proper counting iterators (or ranges), for example, following Thrust or Boost Libraries:
    1counting_iterator <idx_t> first_p(0), last_p(num_particles);
  3. 3.
    The call to the generic algorithm, e.g.:
    1for_each (execution_policy, first_p, last_p, functor);
Step 3 is the one that requires care when performed with a parallel execution policy. Indeed, it requires the handling of possible data races. Currently, we handle them through atomic operations, namely, in the above snippet we substitute m[i] += N * M[k] with atomicAdd ( &m(i), N*M[k]), having re-defined m as a real_t*, and:
  • relying on CUDA implementation of atomicAdd for execution on the GPU,
  • and defining an ad-hoc atomicAdd, based on STL std::atomic_ref, for parallel execution on the CPU.
Atomics cost is not detrimental for overall performance on the GPU, and portability to non-CUDA backends is easily obtained, taking care of defining the ad-hoc atomicAdd for the CPU in such a way to be compliant with CUDA’s API. However, in Section 5 we propose an alternative algorithm which avoids the data-race and thus the need of atomics in P2G like kernels.
@@ 4 Benchmarks
In this section we show the results of 2D numerical simulations of gas dynamics obtained with the described implementation of MPM, with a focus on supersonic gas flow over solid obstacles. A first general evaluation of accuracy is followed by an assessment of the sensitivity of the results to the choice of numerical parameters. We then test the efficiency and performance of the implementation in providing parallel scalability on GPU and its portability to CPUs. The experiments were carried out on NVIDIA A100-40 GPUs and AMD EPYC Rome 7402 CPUs at Leonardo S.p.A.’s in-house HPC facility davinci-1.
@@ 4.1 Test cases
We adopt non-dimensional units as in [41], and we model the fluid as a perfect, bi-atomic gas, with specific heat ratio:
@@
In the far field away from the obstacles, we set fluid density and pressure as:
@@
For time integration, we enforce the time step to be:
@@
In all test cases, initial particles position is sampled by a uniform random distribution in the fluid domain; in the following, we denote with $ppc$ the average number of fluid particles per fluid cell, with $N_{C}$ the total number of cells, and with $N_{P}$ the total number of particles.
@@
First, we consider two cases of supersonic, Mach 3 flow:
@@
  • $\mathbf{v}_{inflow}=(3,0)$
  • $\mathbf{v}_{top}\cdot\mathbf{n}=\mathbf{v}_{bottom}\cdot\mathbf{n}=0$
where $\mathbf{n}$ denotes the unit normal vector. A particles-recycling approach is adopted, whereby particles leaving the domain at the outflow boundary are regenerated in a random cell at the inflow boundary with inflow conditions, keeping the total particles number constant throughout the simulation.
Figure 1: Mach 3 flow past a cylinder: pressure at $t=10\,s$, $N_{C}=10^{6}$, $ppc=16$, $N_{P}=15.6\cdot 10^{6}$ , artificial viscosity as in [56].
In test case (a), as gas particles travel past the cylinder, a shock wave develops at the front and two oblique shocks in the rear. The first shock is reflected by walls and interacts with the back shocks; see Figure 1 and its caption for numerical details.
@@
Figure 2: Mach 3 flow past a step: pressure at $t=12s$, $N_{C}=326.7\cdot 10^{3}$, $ppc=16$, $N_{P}=4.4\cdot 10^{6}$, artificial viscosity as in [56].
The MPM implementation is able to correctly simulate the main flow features, in case of both round shapes and sharp edges and corners. Computed shock-waves angle and flow deflection angle (Figure 1) are found to be in agreement with experimental references [18] as well as with results with other codes, e.g., deal.II, when the same polynomial order is used for basis functions [^5] https://www.dealii.org/current/doxygen/deal.II/step_69.html#Results. The lower level of detail in comparison with some results in the literature ([14, 33, 24]) is likely due to the first order space discretisation used here.
@@
Next, we test the method in simulating flow past an OAT15a aerofoil at Mach 0.73, to assess the ability of the implementation in dealing with more complex-shaped obstacles. The signed-distance function approach, in conjunction with the implemented obstacles treatment (10),(11), naturally fits the aerofoil profile (Figure 3). Since, rather than using an analytic expression, the profile and its normal unit vectors are generated reading from a file containing a set of points, the method proves to be sufficiently general and not bounded to particular geometric shapes.
@@
Figure 4: OAT15a Mach 0.73 flow: velocity streamlines at $\alpha=25^{\circ}$, $t=1s$, $N_{C}=247\cdot 10^{3}$, $ppc=9$, $N_{P}=2.18\cdot 10^{6}$, artificial viscosity as in [56].
These are found to be in line with what is expected from physical considerations: we can see airflow speeding up as it approaches the aerofoil, with an higher velocity magnitude in the top area with respect to the bottom one.
@@
  • $v_{x}(x,y)=v_{\infty}\sin\left(\tfrac{x}{\pi}\right)\cos\left(\tfrac{y}{\pi}\right)$;
  • $v_{y}(x,y)=-v_{\infty}\cos\left(\tfrac{x}{\pi}\right)\sin\left(\tfrac{y}{\pi}\right)$;
  • $p(x,y)=p_{\infty}+\tfrac{\rho_{\infty}}{16}\left(\left(\cos\left(\tfrac{2x}{\pi}\right)+\cos\left(\tfrac{2y}{\pi}\right)\right)\cdot 3-2\right)$.
@@
Figure 5: Taylor-Green vortex: computed velocity at $t=2s$, $N_{C}=3.94\cdot 10^{5}$, $ppc=9$, $N_{P}=3.55\cdot 10^{6}$, artificial viscosity as in [56].

### v2

@@ 3 Portable MPM Implementation
    Data generation: the creation of simulations input is performed separately from code execution. For simple geometries it is handled via ad-hoc C++ scripts; for complex geometries, requiring a careful computation of the signed distance function, an Octave script is used instead, to take advantage of Octave built-in functions. In both cases, since objects are considered at rest, normal unit vectors are computed only once, and boundary cells are marked with a flag corresponding to the boundary condition. Finally, the initial data file is saved in JSON format.
@@
      Equations (12) and (17) are applied to map particles mass, momentum and force to the grid; during force P2GD, artificial viscosity is applied as in (5).
@@
      Velocity and acceleration generalised G2P (16) is applied, and next, velocity is updated. Acceleration itself is not stored, but computed as the result coming from a local G2P of the relevant nodes ratio between force and mass; moreover, G2P kernels are performed on particles index, as shown in the snippet in Fig. 1, since this has proved to provide some acceleration with respect to an approach based on grid nodes index. Following standard MPM [14], we store the G2P velocity in a temporary velocity vector $\mathbf{\overline{V}}$ and don’t use it to update particles velocity; however, we postpone its usage.
@@
      Particles strain rate tensor (velocity gradient) is computed as the G2PD (18) of the ratio between grid nodes momentum and mass, and then particles properties are updated; also in this case, the particle index is used for the G2PD kernel, and artificial viscosity is applied in the update of internal energy density.
@@
      Optionally, a re-ordering algorithm is applied to keep data locality, i.e. to keep contiguous particles in the physical simulation also close in memory. The algorithm is based on an indexing vector built iterating over the structure that stores information about which particles are contained in every grid cell, such that the n-th element of the ordering vector indicates the current sorting particle index which has to be at n-th position after re-ordering.
@@
Algorithm 1 summarises the for loop procedure.
@@
$\qquad m=\mathrm{P2G}(M),\ m\mathbf{v}=\mathrm{P2G}(M\mathbf{V}),\ \mathbf{f}=\mathrm{P2GD}((P+Q)\tfrac{M}{\rho})$
@@
$\qquad\mathbf{V}^{\prime}=\mathbf{V}+\Delta t\mathbf{A}$
@@
$\qquad E^{\prime}=E+\Delta t\dfrac{-(P+Q)}{\rho}\nabla\cdot\mathbf{V}$
@@
This algorithm is of special interest for acceleration on GPUs since it can be written in terms of the four kernels (15)-(18), which present interesting parallelisation potentialities. Grid-to-Particles operations are intrinsically Single Instruction Multiple Data (SIMD) parallelisable, since data is projected from grid nodes to particles, and thus particles are independent from one another during writing stage. Particles-to-Grid kernels present the same potentialities, however, if they are implemented strictly following their analytical definition, they present a data race: different particles need access in both read and write mode to the same grid nodes.
In this work, we chose to patch the Particles-to-Grid and Grid-to-Particles kernels written in the external quadgrid library to adapt them to GPU execution via Thrust rather than writing ad hoc versions in the code, at the expense of some further optimisation that could have been added by exploiting the peculiarities of our test cases, to arrive at more general and portable conclusions. In this way, there is only one computational kernel per operation (15)-(18). In any case, preliminary tests have shown that applying a kernel-fusion approach [44] where possible wouldn’t have affected the execution time order of magnitude.
To better illustrate the concepts introduced in the last paragraph, we sketch and discuss below the implementation and use of the two main kernels in the algorithm, namely the Grid-to-Particle (G2P) operation (13) and the Particle-to-Grid operation (12). A simplified sketch of the implementation of the former is given in Fig. 1, while the latter is described in Fig. 4; the reported code corresponds roughly to the actual implementation available in the quadgrid library except for a few details that are left out only to avoid distraction as they are not interesting for the current discussion. As already noted, the G2P operation is easily implemented via a loop over all particles and lends itself quite naturally to SIMD parallelisation as all write operations are independent and may be carried out concurrently. The actual looping is performed via the for_each method, which can be used to perform any operation on any number of variables associated to a particle by using a counting iterator and a suitably constructed callable object as shown in Fig. 3 and Fig. 2. The P2G operation is implemented along lines very similar to those followed for G2P. Specifically, as shown in Fig. 4 the looping is done over the list of particles just as for G2P, which is made possible by the index array ptcl_to_grid that lists, for each particle, the grid cell within which it is located. Efficiently building the latter data structure heavily depends on the use of structured grid. The main difference between P2G and G2P consists in the need, in the former, to work around a data race that is created when threads working on different particles try to write data related to the same neighbouring cell vertex. The simplest approach, followed in this work, is to make the write operations atomic by using atomicAdd (as shown for sake of example in Fig. 4), which is readily available in CUDA but can also be re-implemented in standard C++ based on the std::atomic_ref class template introduced in the C++20 standard. Although our tests show that the use of atomics does not seem to lead to much of a performance degradation on GPUs, we discuss in a later section a possible alternative based on data reordering and mesh colouring that can be used to avoid atomics.
1 template<class callable_t>
2 G2P (idx_t iptcl, callable_t *grid_variable,
3 ptcl_variable_t *ptcl_variable) {
4
5 idx_t icell = ptcl_to_grid[iptcl];
6 real_t x = ptcl_x[iptcl];
7 real_t y = ptcl_y[iptcl];
8
9 for (idx_t inode = 0; inode < 4; ++inode) {
10 real_t N = shape_function (x, y, inode, icell);
11 idx_t ginode = local_to_global (icell, inode);
12
13 ptcl_variable[iptcl] +=
14 N * (*grid_variable)(ginode);
@@
16
17 };
Figure 1: Example template function to demonstrate the approach used to implement the G2P kernel. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In particular note that in this simplified snippet (and some of the following) we use, for sake of simplicity of presentation, standard pointers, while the actual implementation uses some form of iterator.
1 struct x_velocity_t {
2 grid_data_t *gdata = &grid_data;
3 real_t operator()(idx_t ginode) {
4 real_t vx = gdata -> get (ginode, "x_momentum");
5 real_t m = gdata -> get (ginode, "mass");
6 return vx / m;
7 }
8 } *x_velocity;
9
10 x_velocity_g2p =
11 [x_velocity, ptcl_x_velocity] (idx_t iptcl) {
12 G2P (iptcl, x_velocity, ptcl_x_velocity);
13 };
Figure 2: Instantiation of a callable object that implements grid to particle interpolation of the $x$–component of the velocity. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In the code above, the purpose of the callable object pointed to by x_velocity is to compute on–the–fly a function of the state variables at a grid vertex, so that its value can be interpolated at the particle locations without requiring extra storage or data transfers.
1 counting_iterator<idx_t> first_p(0);
2 ounting_iterator<idx_t> last_p(num_particles);
3 for_each (execution_policy,
4 first_p, last_p, x_velocity);
Figure 3: Example of performing an actual G2P operation by iterating over all particles by means of the for_each method and counting iterators.
1 template<class callable_t>
2 P2G (idx_t iptcl, ptcl_variable_t *grid_variable,
3 callable_t *ptcl_variable,) {
4
5 idx_t icell = ptcl_to_grid[iptcl];
6 real_t x = ptcl_x[iptcl];
7 real_t y = ptcl_y[iptcl];
8
@@ 2 P2G (idx_t iptcl, ptcl_variable_t *grid_variable,
10 real_t N = shape_function (x, y, inode, icell);
11 idx_t ginode = local_to_global (icell, inode);
12
13 atomicAdd (grid_variable + iptcl,
14 N * (*ptcl_variable)(ginode));
@@
16
17 };
Figure 4: Example template function to demonstrate the approach used to implement the G2P kernel. This is not the actual code used but a conceptualised, reduced example to explain the basic idea. In particular note that in this simplified snippet (and some of the following) we use, for sake of simplicity of presentation, standard pointers, while the actual implementation uses some form of iterator.
@@ 4 Benchmarks
In this section we show the results of 2D numerical simulations of gas dynamics obtained with the described implementation of MPM, with a focus on transonic and supersonic gas flow over solid obstacles. A first general evaluation of accuracy in subsection 4.1 is followed by an assessment of the sensitivity of the results to the choice of numerical parameters in subsection 4.2; in particular, the effects of numerical viscosity components on mass density are shown for a test case with analytical solution in subsection 4.2.
In subsection 4.3 we then test the efficiency and performance of the implementation in providing parallel scalability on NVIDIA A100 GPUs. Lastly, in subsection 4.4, we assess the portability of the developed code to different CPUs and GPUs architecture, which is our main claim.
The experiments were carried out on NVIDIA A100 SXM4 40 GB GPUs, Intel Xeon Platinum 8260 and AMD EPYC Rome 7402 CPUs at Leonardo S.p.A.’s in-house HPC facility davinci-1, and on AMD EPYC 7313, AMD EPYC 7413 CPUs and AMD Radeon Instinct MI210 GPUs at MOX Laboratory for Modeling and Scientific Computing at Politecnico di Milano’s Mathematics Department.
All tests are performed in double precision, i.e. using FP64. Depending on the architecture, better performance could be achieved with lower precision floating points; for example, the NVIDIA GPUs here employed, which date back to 2020, should gain a 2x factor in computational time when working with FP32. However, despite the code supports different choices, we have focused only on doubles, since floats could lead to numerical difficulties in some test cases, and already the slightly more recent AMD MI 210 GPUs (release date December 2021) are able to handle double and single precision floats without any performance degradation.
@@ 4.1 Test cases
In the following numerical tests, we adopt non-dimensional units as in [47], and we model the fluid as a perfect, bi-atomic gas, with specific heat ratio:
@@
Moreover, except in the Sod Shock Tube, in the far field away from any obstacle, we set fluid density and pressure as:
@@
For time integration, in our code we enforce the time step to be:
@@
In all test cases except in the Sod Shock Tube, initial particles position is sampled by a uniform random distribution in the fluid domain. Indeed, a sufficient randomness of particles position, especially in particles recycling from outlet to inlet, is fundamental to obtain stable and physically meaningful simulations, in agreement with results in the literature on Particle-In-Cell methods in plasma physics. On longer runs (i.e. above tens of physical seconds) a symplectic time integration scheme might be useful [34].
In the following, we denote with $ppc$ the average number of fluid particles per fluid cell, with $N_{C}$ the total number of cells, and with $N_{P}$ the total number of particles.
Sod Shock Tube
In the Sod Shock Tube, two compressible fluids, initially at rest, are divided by a diaphragm. The first fluid has $\rho_{1}=1$ and $p_{1}=1$, the second $\rho_{2}=0.125$ and $p_{2}=0.1$. At time $t=0$ the diaphragm is removed, and the system evolution is studied; the following phenomena are expected:
  • the two initial zones shrink progressively towards domain borders;
  • a rarefaction wave propagates toward the high density zone;
  • in the position where the diaphragm was, which now is moving towards low densities, a contact discontinuity develops;
  • a shock wave propagates towards low density zones.
In order to reproduce the 1D problem, we introduce a mesh of $1\times 100$ cells; moreover, to be able to compare the MPM solution with the reference one, we also set initial particle position as in [65, 64].
In Figure 5 we show the final particles position, together with a comparison of the computed solution, its linear interpolation on the grid, and the exact solution computed following the procedure shown in [56]. Results are in very good agreement with the MPM literature [65, 19].
Figure 5: Sod Shock Tube: comparison of MPM solution computed on particles, interpolated on the grid, and exact solution computed as in [56]
@@
We then consider two cases of supersonic, Mach 3 flow:
@@
$\begin{cases}\mathbf{v}_{inflow}=(3,0)\\ \mathbf{v}_{top}\cdot\mathbf{n}=\mathbf{v}_{bottom}\cdot\mathbf{n}=0\end{cases}$
where $\mathbf{n}$ denotes the unit normal vector. A particles-recycling approach is adopted, whereby particles leaving the domain at the outflow boundary are regenerated in a random cell at the inflow boundary with inflow conditions, keeping the total number of particles constant throughout the simulation. In test case (a), as gas particles travel past the cylinder, a shock wave develops at the front and two oblique shocks in the rear. The first shock is reflected by walls and interacts with the back shocks; see Figure 6 and its caption for numerical details.
Figure 6: Mach 3 flow past a cylinder: pressure at $t=10\,s$, $N_{C}=10^{6}$, $ppc=16$, $N_{P}=15.6\cdot 10^{6}$
Figure 7: Mach 3 flow past a step: pressure at $t=12s$, $N_{C}=326.7\cdot 10^{3}$, $ppc=16$, $N_{P}=4.4\cdot 10^{6}$
@@
The MPM implementation is able to correctly simulate the main flow features, in case of both round shapes and sharp edges and corners. Computed shock-waves angle and flow deflection angle (Figure 6) are found to be in agreement with experimental references [20] as well as with results with other codes, e.g., deal.II, when the same polynomial order is used for basis functions. In particular, the cylinder test case is simulated in the deal.II documentation [39] [^7] https://www.dealii.org/current/doxygen/deal.II/step_69.html#Results. With our implementation, we are able to reproduce the main features of the front shock starting from a very low resolution, with details, especially in the rear, getting better as resolution increases; more details are shown in subsection 4.2. Overall, the results are in agreement with [39], while the lower level of detail in comparison with some results in the literature ([16, 38, 27]) is likely due to the first order space discretisation used here.
@@
Next, we test the method in simulating compressible, transonic flow past an OAT15a aerofoil at Mach 0.73, to assess the ability of the implementation in dealing with more complex-shaped obstacles. The signed-distance function approach, in conjunction with the implemented obstacles treatment (19),(20), naturally fits the aerofoil profile (Figure 8). Since, rather than using an analytic expression, the profile and its normal unit vectors are generated reading from a file containing a set of points, the method proves to be sufficiently general and not bounded to particular geometric shapes.
@@
These are found to be in line with what is expected from physical considerations: we can see airflow speeding up as it approaches the aerofoil, with a higher velocity magnitude in the top area with respect to the bottom one.
Figure 9: OAT15a Mach 0.73 flow: velocity streamlines at $\alpha=25^{\circ}$, $t=1s$, $N_{C}=247\cdot 10^{3}$, $ppc=9$, $N_{P}=2.18\cdot 10^{6}$.
@@
$\begin{cases}v_{x}(x,y)=v_{\infty}\sin\left(\tfrac{x}{\pi}\right)\cos\left(\tfrac{y}{\pi}\right)\\ v_{y}(x,y)=-v_{\infty}\cos\left(\tfrac{x}{\pi}\right)\sin\left(\tfrac{y}{\pi}\right)\\ p(x,y)=p_{\infty}+\tfrac{\rho_{\infty}}{16}\left(\left(\cos\left(\tfrac{2x}{\pi}\right)+\cos\left(\tfrac{2y}{\pi}\right)\right)\cdot 3-2\right).\end{cases}$
@@
Figure 10: Taylor-Green vortex: computed velocity at $t=2s$, $N_{C}=3.94\cdot 10^{5}$, $ppc=9$, $N_{P}=3.55\cdot 10^{6}$